How long will it take a butterfly going 11.5 mph to travel 2500 miles with no rest
Answer by TimothyLamb(4379)

v = d / t
t = d / v
t = 2500 / 11.5 = 217.4 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php